# LLMs Reproduce Stereotypes of Sexual and Gender Minorities

### v1

@@ Abstract
A large body of research has found substantial gender bias in NLP systems. Most of this research takes a binary, essentialist view of gender: limiting its variation to the categories men and women, conflating gender with sex, and ignoring different sexual identities. But gender and sexuality exist on a spectrum, so in this paper we study the biases of large language models (LLMs) towards sexual and gender minorities beyond binary categories. Grounding our study in a widely used psychological framework—the Stereotype Content Model—we demonstrate that English-language survey questions about social perceptions elicit more negative stereotypes of sexual and gender minorities from LLMs, just as they do from humans. We then extend this framework to a more realistic use case: text generation. Our analysis shows that LLMs generate stereotyped representations of sexual and gender minorities in this setting, raising concerns about their capacity to amplify representational harms in creative writing, a widely promoted use case.
@@ 1 Introduction
Research has established that a host of biases based on characteristics such as gender, race, sexuality, and nationality are present in LLMs (Navigli et al., 2023), and in NLP more broadly. A substantial proportion of this research has focused specifically on gender, but recent surveys (Stanczak and Augenstein, 2021; Devinney et al., 2022) have found that most papers in this area take an oversimplified view of gender, typically treating it as binary (by considering only the genders women and men) and essentialist (conflating gender with physical characteristics, and often implicitly with sexuality).
@@
This paper expands recent efforts to extend the study of gender bias in LLMs beyond these oversimplifications (Dhingra et al., 2023). We look specifically at bias towards a spectrum of gender and sexual minorities, and we focus on creative text generation, a use case that has been widely promoted in the marketing materials of LLM providers, including the providers of ChatGPT [^1] https://openai.com/chatgpt/use-cases/writing-with-ai/, Gemini [^2] https://sites.google.com/view/gemini-workspace-ai, and LLaMA [^3] https://ai.meta.com/blog/meta-llama-3/, the LLMs that we study in this paper.
@@
Following the recommendations of Blodgett et al. (2020), we aim to connect bias in language models to possible harms, and following the recommendations of Goldfarb-Tarrant et al. (2023), we aim to ground our operationalization of bias in established models of measurement. One harm that can result from text generation is representational harm (Crawford, 2017), the perpetuation and amplification of harmful stereotypes about a social group, which can in turn reinforce other harmful behaviors towards members of that group. To assess the risk of representational harm, we need an operational definition of stereotype. For this purpose, we employ the Stereotype Content Model (SCM; Fiske et al., 2002), a widely used framework from social psychology research (Section 2), which has recently been adopted for research on bias in NLP systems (e.g. Ungless et al., 2022), including research on bias in LLMs (Jeoung et al., 2023; Salinas et al., 2023). Highly influential in social psychology research, the SCM models stereotypes of groups as differentiated along axes of Warmth and Competence. Importantly, there is evidence that behavior towards social groups is correlated with perceptions of stereotype on these axes (Cuddy et al., 2007), thus linking representational harm to further harms.
@@
To assess whether LLMs reproduce stereotypes of sexual and gender minorities, we first use the methodology of the SCM (Section 3) to establish a correlation between the responses of human participants and LLMs. Extending recent work of Jeoung et al. (2023), we prompt both human participants and LLMs with survey questions about social perceptions of sexual and gender minorities, finding that LLMs do indeed reflect the behavior of human participants both quantitatively and qualitatively (Section 4). These results are not surprising, but the survey task is artificial, and not representative of real LLM use cases. So, we extend our analysis to text generation by mapping generated words onto the SCM axes of Warmth and Competence using semantic similarity.
@@
Our results replicate previous findings about the binary categories of women and men, with women strongly rated for Warmth and men for Competence. They also show that LLMs produce more negative representations of bisexual and nonbinary people, with descriptions focused on lived hardships. Some differences are apparent in the LLMs, with Gemini the most divergent of the models.
@@ 2 Background
The Stereotype Content Model (SCM; Fiske et al., 2002) is a general framework which theorizes that many culturally-specific stereotypes can often be reduced to a pair of dimensions, Warmth and Competence, discussed in more detail below. The SCM is well-established and many of its details have been validated through multiple studies (e.g. Fiske et al., 2002; Fiske, 2018; Cuddy et al., 2008; Nicolas et al., 2021). Though originating in the United States, it has been reproduced in several cultural contexts, consistently showing that outgroups are perceived more negatively on one or both axes Cuddy et al. (2009).
@@
The SCM does not conceptualize stereotypes solely in terms of negative or positive views of a given group. Rather, it theorizes that stereotypes can be reduced to perceptions based on Warmth and Competence Fiske et al. (2002). Given perceptions along these two axes, groups can be mapped into four quadrants, each defined by whether it contains low or high values along each axis. Cuddy et al. (2007) showed that public perceptions associated with these different quadrants are statistically linked to both emotions and behaviors: for example, the Low Warmth / Low Competence quadrant is associated with the emotion of contempt (Figure 1), and members of groups in this quadrant are likely to be subject to the behavior of harm—both active harm, like harassment, and passive harm, like neglect. In this way, the SCM links the representational harms associated with stereotypes to further real world harms.
@@
Figure 1: The Bias Map (reproduced from Cuddy et al., 2007) illustrates how different social stereotypes of groups on the Warmth and Competence axes relate to emotions expressed towards those groups (gray arrows) and behaviors towards those groups (black arrows). Cuddy et al. (2007) found that stereotypes of a group are empirically correlated with both emotion and behavior towards that group in the directions that the bias map hypothesizes.
The SCM has been used to study stereotypes in NLP systems for several years. Mieczkowski et al. (2019) used it to measure human perceptions of robots. More relevant to us, it has been used to study biases in masked language models related to disability and mental illness (Herold et al., 2022; Mina et al., 2024)), and biases in LLMs (Salinas et al., 2023; Jeoung et al., 2023). We take particular inspiration from StereoMap (Jeoung et al., 2023), a prompting framework for analyzing LLMs’ perceptions of different social groups. Jeoung et al. (2023) use StereoMap to establish a correlation between LLM behavior and Fiske et al. (2002)’s social psychological findings for many social groups along the axes of the SCM, suggesting that the SCM is a valid instrument for these measurements. Our goal is to extend their work by considering a new set of social groups, sexual and gender minorities, and by extending it to the more realistic setting of text generation.
@@ 3 Methodology
Our approach requires a way to operationalize the set of groups that we study and the concepts of Warmth and Competence, in addition to a set of language models.
@@ 3.1 Group and attribute terminology
For gender and sexuality groups that have been studied in previous research with the SCM—women, men, and gay men—we use the same terminology. Since we also consider groups that were not in those studies, we ran pilot tests with LLMs comparing results based on similar terms (e.g., “heterosexual” and “straight”; “nonbinary” and “gender fluid”). We found that similar terms returned similar results, so we used the most common terms. We acknowledge that the set of identities used in our experiments (Table 1) does not constitute an exhaustive list, but keeping the list focused was critical for surveying human participants without causing fatigue.
Our survey method also requires Warmth and Competence terms. For these, we combine word lists from multiple studies employing the SCM (Fiske et al., 2002; Cuddy et al., 2008; Jeoung et al., 2023), producing a more comprehensive list of eleven terms for each axis (Table 1). For surveys of human participants (Section 4.3), we use only the Fiske et al. (2002) subset in order to prevent survey fatigue. All words in these lists are positive, following Fiske et al. (2002). This is because rating groups on these attributes is part of the survey method, and we confirmed in pilot experiments that LLMs generally refuse to explicitly rate social groups against negative attributes.
@@ 3.2 Models
We tested three representative models that were in widespread use at the time of our work: GPT 3.5-turbo, Gemini-1.5-flash, and LLaMA 2-7b-chat-hf. GPT and Gemini were accessed via API, whereas the smaller, open-source LLaMA was run locally. Following pilot experiments, we used a temperature of 0.9 for GPT and Gemini, which were relatively insensitive to this parameter. For LLaMA, we used a temperature of 0.6 and top-$p$ of 0.9, finding that higher temperatures produced output unrelated to the prompt or in other languages.
The LLMs in our experiments contain safety settings which are intended to block harassment, hate speech, sexual content, and dangerous language. Although our experiments don’t contain such material, they do explicitly mention social groups that are frequent targets of such material. In pilot experiments, we found that these safety settings resulted in frequent refusals to produce the requested output. For example, Gemini blocked approximately one third of our prompts in both survey and text generation experiments. An analysis showed negligible difference in output for unblocked prompts, so we turned off the safety setting for the main experiments reported below.
@@
4 SCM Survey of Humans and Prompting of LLMs
We first want to understand whether LLMs behave similarly to humans on an established task: the SCM survey, which assesses societal perceptions of groups on the Warmth-Competence axes. This follows StereoMap (Jeoung et al., 2023), which compares pre-existing human survey results from Fiske et al. (2002) to LLM behavior on a set of prompts inspired by the original survey. But the survey by Fiske et al. (2002) only includes women, men, and gay men; it does not include nonbinary people, lesbians, bisexuals, or heterosexuals. In order for us to compare LLM biases to societal biases towards these groups, we conducted a new survey of human participants. We ask both human participants and LLMs to answer the same types of questions, which probe societal perceptions of groups on the Warmth and Competence axes in three ways:
@@ 4.1 LLM Prompt Design
LLMs were tested using a set of prompts inspired by StereoMap Jeoung et al. (2023), themselves adapted from the survey questions of Fiske et al. (2002), mirroring the three tasks that we gave to human participants. Note that Jeoung et al. (2023) did not include a selection task. We include one here to better align with previous research on stereotypes (e.g. Cao et al., 2022; Shrawgi et al., 2024; Nadeem et al., 2021). Using multiple types of survey question also gives us multiple views on the behavior of each LLM, and is more robust to varying rates of refusals.
@@
Table 2: Sample LLM prompts, showing variables in teal. Following StereoMap (Jeoung et al., 2023), the request for models to explain their reasoning, as seen in the first prompt, is based on the rationale of Chain Of Thought (CoT) prompting Wei et al. (2022), since this often yields improvements to performance on reasoning tasks .
@@ 4.2 Associating Elicited Words to SCM Axes
Since elicited words are not necessarily from our predefined lists, we will need a way to determine which of them represent concepts of Warmth and Competence, in order to relate the outcome to other survey questions. To do this, we use a dictionary created by Nicolas et al. (2021) as a psychological measurement instrument. It associates a large number of terms with several widely-studied stereotype dimensions, partly inspired by the SCM concepts of Warmth and Competence, and has been tested for internal consistency and validity with respect to human judgment, as well as other psychological inventories used to measure these dimensions. For purposes of our analysis, we associate their categories of Morality and Sociability with Warmth, and their categories of Agency and Ability with Competence. Words in these categories account for 45% of the observed word types in our elicited data. For the remaining words, we compute the cosine similarity of their OpenAI word vectors to the average word vector of our Warmth and Competence keywords (Table 1). We assume that those words with a similarity greater than 0.55 represent the associated concept, since this threshold gave us the highest agreement for words in the Nicolas et al. (2021) inventory.
4.3 Human Responses to SCM Survey
@@
Using responses from our survey of human participants, we first establish what the social stereotypes of these groups are. We recruited participants using the platform Prolific, filtering for English fluency. Ninety-seven participants were each asked to answer twenty-one questions—one of each type, for each of the seven groups in Table 1. They were compensated with an amount above the national minimum wage in the country where we conducted our research.
In order to understand our participants, we included two demographic questions about age range and gender identity, but retained no further identifying information. Most participants were between 16 and 35 (79%), with 14% between 36 and 45, with similar numbers of women and men (54% and 43%, respectively). 6% of participants were over the age of 45 and 2% of participants identified as nonbinary.
Human Rating
@@
Figure 2 summarizes the results of the rating question by plotting the Warmth and Competence ratings for each group in two dimensions, clearly showing that perceptions of each group do indeed differ. Heterosexuals are rated highly in both axes, while nonbinary people are rated low in both. Gay, lesbian, and bisexual groups are clustered between the outliers. The results for women, men, and gay men are consistent with those of Fiske et al. (2002), who did not include the other groups in our suvey. In particular, women and men rate most highly in Warmth and Competence, respectively.
222233334444222233334444bisexualgayheterosexuallesbianmennonbinarywomenAverage Competence RatingAverage Warmth Rating
Figure 2: Average human rating of Warmth and Competence for each group. In principle, ratings can range from 1 to 5, but in practice, they fell between 2 and 4, so we show only that range. All other results tables and figures in this paper sort the groups by ascending order of their human rating for Competence, in order to facilitate comparison with these results.
Human Selection
The second question of our survey investigates societal perceptions by asking participants to select a single term from a list of twenty. Ten of these terms are a subset of the Warmth-Competence key terms (Table 3) and the remaining ten are the inverse of each positive word, such as ‘cold’ for ‘warm’. The results (Figure 3) are broadly consistent with the rating results and with previous studies: women are rated highest for Warmth; men rated highest for Competence; nonbinary people rated most negatively; heterosexuals rated most positively. The main notable difference from the Rating task is a slight preference for Competence terms selected for gay men, contrasting with a slight preference for Warmth in the rating task.
womenmennonbinarygaylesbianbisexualheterosexual2020-20- 200020202020404040406060606080808080% of responses (“not" terms in negative)CompetentNot CompetentWarmNot Warm
@@
Figure 3: Percentages for human selection of Warmth and Competence terms, and their inverses, for each group. Since this is a selection task, the sum of all four percentages is 100% for each group. To make comparisons of Warmth and Competence easier, we show them side by side, and to make comparisons of overall positive and negative terms easier, we show the negative terms as negative values on the vertical axis.
Human Elicitation
The five most frequently elicited keywords per group are summarized in Table 3, while rates of Warmth or Competence-related words amongst all keywords are summarized in Figure 4. In both cases the method in Section 4.2 is used to identify words related to Warmth and Competence. We again see similar patterns to the other survey question results: terms elicited for women tend to associate with Warmth (e.g. “nurturing”), while those for men associate with Competence (e.g. “leader”). Heterosexuals elicit strong associations for both Warmth and Competence, and, more qualitatively, with normalcy (e.g. “normal” and “natural”). In contrast, nonbinary and bisexual people elicit more negative terms, including words relating to confusion (e.g. “confused”, “lost”, “indecisive”). Notably, all sexual and gender minorities elicited fewer Warmth and Competence terms than women, men, or heterosexual groups.
@@
confused emotional confused flamboyant masculine normal strong
@@
Table 3: Most frequent keywords for each group obtained by human elicitation. Words in orange are associated with Warmth and words in blue with Competence using the method in Section 4.2.
womenmennonbinarygaylesbianbisexualheterosexual10101010202020203030303040404040% of responsesCompetenceWarmth
Figure 4: Percentage of Warmth and Competence terms obtained by human elicitation, using the method of Section 4.2 to associate words to these concepts. Since terms unrelated to either concept occur in the data, percentages do not sum to 100.
Critically, the different types of survey question yield consistent results: they recapitulate SCM findings about stereotypes of men, women, and gay men (Fiske et al., 2002), and repeatedly elicit a perception of more negative stereotypes of sexual and gender minorities, most strongly of nonbinary and bisexual people.
4.4 LLM Responses to SCM Prompts
LLM Rating
Results for all three models on the rating task are plotted in Figure 5. Though there is some difference in the magnitude of the ratings, the general pattern for both GPT and LLaMA models is similar to the one that we observe in the human survey. In particular, women are rated most highly for Warmth and men for Competence, with heterosexual people rated highly on both axes and nonbinary people rated lowest for Competence and relatively low for Warmth. Gay men, lesbians, and bisexual people cluster between the bounds set by the other groups. Gemini behaves differently, exhibiting an inverse relationship between Warmth and Competence, though women and men still bound the ratings with highest values for Warmth and Competence, respectively.
3333444433334444GPTbisexualgayheterosexuallesbianmennonbinarywomenAverage Competence RatingAverage Warmth Rating
3333444433334444GeminibisexualgayheterosexuallesbianmennonbinarywomenAverage Competence Rating
3333444433334444LLaMAbisexualgayheterosexuallesbianmennonbinarywomenAverage Competence Rating
Figure 5: Average LLM ratings of Warmth and Competence for each group (cf. Figure 2). In principle, ratings can range from 1 to 5, but in practice, they only fell in a narrower range that we show here. Note that the LLaMA ratings for nonbinary and bisexual people are nearly identical, so are difficult to distinguish in the visualization.
LLM Selection
In the selection task, we report results for GPT and Gemini in Figure 6, since LLaMA had a very high rate of refusal for these prompts. GPT is somewhat consistent with human survey results. Gemini tends to skew strongly towards either Warmth or Competence for each group, though this skewed response is internally consistent with its rating results. Additionally, it prefers Competence for three groups that humans rated most highly for Competence.
00252525255050505075757575100100100100GPTof responsesCompetenceWarmth
Figure 6: Percentage of Warmth and Competence terms obtained by LLM selection. Because this is a selection task, percentages necessarily sum to 100. LLaMA omitted due to high refusal rate.
LLM Elicitation
LLMs produced associations with Warmth and Competence that are relatively consistent with previous results both quantitatively (Figure 7) and qualitatively (Table 4). For GPT and LLaMA, rates of Warmth and Competence-related keywords tend to follow the relative patterns of human ratings, with GPT following them quite closely. Nonbinary people and bisexuals receive particularly low rates of terms coded for either Warmth or Competence. Gemini remains an outlier: results for each group often contain high rates of either Warmth or Competence, but rarely both. Note that this skew differs from the one observed in the selection task, where, for nonbinary people, bisexuals, and heterosexuals, the preferred category differs between tasks.
Qualitatively, frequently elicited keywords are distinct for each group, and several are common to the human elicitation, repeated by multiple models, or both. Keywords for the minority groups include “courageous”, “brave”, “strength” and “resilience”, which are coded for Competence but also allude to historical discrimination (Table 4). Creativity is also a theme across all sexual and gender minority groups.
00252525255050505075757575100100100100GPTof termsCompetenceWarmth
00252525255050505075757575100100100100Geminiof terms
Figure 7: Percentage of Warmth and Competence terms obtained by LLM elicitation (cf. Figure 4). Since terms unrelated to these concepts occur in the data, percentages do not sum to 100.
Nonbinary Women Bisexual Gay Lesbian Heterosexual Men

### v2

@@ Abstract
A large body of research has found substantial gender bias in NLP systems. Most of this research takes a binary, essentialist view of gender: limiting its variation to the categories men and women, conflating gender with sex, and ignoring different sexual identities. But gender and sexuality exist on a spectrum, so in this paper we study the biases of large language models (LLMs) towards sexual and gender minorities beyond binary categories. Grounding our study in a widely used social psychology model—the Stereotype Content Model—we demonstrate that English-language survey questions about social perceptions elicit more negative stereotypes of sexual and gender minorities from both humans and LLMs. We then extend this framework to a more realistic use case: text generation. Our analysis shows that LLMs generate stereotyped representations of sexual and gender minorities in this setting, showing that they amplify representational harms in creative writing, a widely advertised use for LLMs.
LLMs Reproduce Stereotypes of Sexual and Gender Minorities
Ruby Ostrowthanks: Work completed while at the University of Edinburgh ruby.a.ostrow@gmail.com                        Adam Lopez University of Edinburgh alopez@ed.ac.uk
@@ 1 Introduction
Research has established that a host of biases conditioned on gender, race, sexuality, and nationality are present in LLMs (Navigli et al., 2023), and in NLP more broadly. Most of this research has focused specifically on gender, but recent surveys (Stanczak and Augenstein, 2021; Devinney et al., 2022) have found that this research takes an oversimplified view of gender, treating it as binary (by considering only the genders men and women) and essentialist (conflating gender with physical characteristics, and often implicitly with sexuality).
@@
This paper expands on efforts to study gender bias in LLMs beyond these oversimplifications (Dhingra et al., 2023). We aim to measure bias towards gender and sexual minorities in creative text generation, a use case that has been widely advertised by LLM providers, including the providers of ChatGPT [^1] https://openai.com/chatgpt/use-cases/writing-with-ai/, Gemini [^2] https://sites.google.com/view/gemini-workspace-ai, and LLaMA [^3] https://ai.meta.com/blog/meta-llama-3/, the LLMs that we study in this paper.
@@
Following Blodgett et al. (2020), we aim to connect bias to possible harms, and following Goldfarb-Tarrant et al. (2023), we ground our operationalization of bias in an established model of measurement. One harm that can result from text generation is representational harm (Crawford, 2017) from perpetuating and amplifying negative stereotypes about a social group, which can reinforce harmful behaviors towards members of that group. To measure representational harm, we need an operational definition of stereotype. For this purpose, we employ the Stereotype Content Model (SCM; Fiske et al., 2002), a widely used framework from social psychology (Section 2), which has previously been used to measure bias in NLP (e.g. Ungless et al., 2022) and in LLMs (Jeoung et al., 2023; Salinas et al., 2023). Highly influential in social psychology research, the SCM models stereotypes of groups as differentiated along axes of Warmth and Competence. Importantly, there is evidence that behavior towards social groups is correlated with perceptions of stereotype on these axes (Cuddy et al., 2007), thereby linking representational harm to further harms.
@@
To assess whether LLMs reproduce stereotypes of sexual and gender minorities, we first use the methodology of the SCM (Section 3) to establish a correlation between the responses of human participants and LLMs. We prompt both human participants and LLMs with survey questions about social perceptions of sexual and gender minorities, finding that LLMs do indeed reflect the behavior of human participants both quantitatively and qualitatively (Section 4). These results are not surprising, but the survey task is artificial, and not representative of real LLM use cases. So, we extend our analysis to text generation by mapping generated words onto the SCM axes of Warmth and Competence using semantic similarity (Section 5).
@@
Our results show that LLMs produce more negative representations of bisexual and nonbinary people, with descriptions focused on lived hardships. Some differences are apparent in the LLMs, with Gemini the most divergent of the models. We strongly advise caution in using these tools to generate text about demographic groups, since they demonstrably reproduce stereotypes.
@@ 2 Background
The Stereotype Content Model (SCM; Fiske et al., 2002) theorizes that many culturally-specific stereotypes can be reduced to a pair of dimensions, Warmth and Competence, discussed in more detail below. The SCM is well-established and its details have been validated through multiple studies (e.g. Fiske et al., 2002; Fiske, 2018; Cuddy et al., 2008; Nicolas et al., 2021). Though originating in the United States, it has been reproduced in several cultural contexts, consistently showing that outgroups are perceived more negatively on one or both axes Cuddy et al. (2009).
@@
The SCM does not conceptualize stereotypes as negative or positive views of a social group. Instead, it theorizes that stereotypes can be reduced to perceptions of Warmth and Competence Fiske et al. (2002). Given perceptions on these axes, groups can be mapped into four quadrants, each defined by low or high values along each axis. Cuddy et al. (2007) showed that perceptions associated with these different quadrants are statistically linked to both emotions and behaviors. For example, the Low Warmth / Low Competence quadrant is associated with the emotion of contempt (Figure 1), and social groups in this quadrant are more often the target of harm—both active harm, like harassment, and passive harm, like neglect. Hence, the SCM links the representational harms of stereotypes with real world harms.
@@
Figure 1: The Bias Map (reproduced from Cuddy et al., 2007) illustrates how different social stereotypes of groups on the Warmth and Competence axes relate to emotions expressed towards those groups (gray arrows) and behaviors towards those groups (black arrows). Cuddy et al. (2007) found that stereotypes of a group are correlated with both emotion and behavior towards that group in these directions.
The SCM has been used to study stereotypes in NLP for several years, in masked language models (e.g. Herold et al., 2022; Mina et al., 2024)) and LLMs (Salinas et al., 2023; Jeoung et al., 2023). We take inspiration from StereoMap (Jeoung et al., 2023) which uses SCM prompts to analyse LLM responses about different social groups. StereoMap established a correlation between LLM behavior and Fiske et al. (2002)’s human survey, validating it as a measurement instrument.
Taking StereoMap’s theoretical validation of the SCM for measuring stereotypes in LLMs as a starting point, we have three goals. First, we focus on sexual and gender minorities, offering much richer quantitative and qualitative evidence for stereotypes of these groups than Jeoung et al. (2023). [^4] Jeoung et al. (2023) include all of the groups in our analysis except for nonbinary people. They conduct a broad statistical analyses over nearly 100 social groups of different overlapping types (e.g. age, social status, job, gender, etc.), focused on validating the SCM as a measure for stereotypes in LLMs; we take their validation as given. In contrast, we deeply analyze stereotypes of previously ignored social groups. Second, we verify that LLM stereotypes resemble those of humans by conducting a new survey with humans. [^5] The human survey results to which Jeoung et al. (2023) compare (Fiske et al., 2002; Cuddy et al., 2007) include only 23 groups, and do not include nonbinary, lesbian, bisexual, or heterosexual groups. Finally, we extend the SCM methodology to analyze text generation.
@@ 3 Methodology
We need a way to operationalize a set of social groups and the concepts of Warmth and Competence; and a set of LLMs.
@@ 3.1 Group and attribute terminology
For gender and sexuality groups that have been studied in previous SCM research—women, men, and gay men—we use the same terminology. Since we consider groups that were not in those studies, we ran pilot tests with LLMs using similar terms for the same social group (e.g., heterosexual and straight; nonbinary and gender fluid). Since similar terms returned similar results, we settled on popular terms. We acknowledge that the list of identities in our experiments (Table 1) is not exhaustive, but keeping it brief was critical for surveying human participants without causing fatigue.
For Warmth and Competence terms, we combine word lists from multiple SCM studies (Fiske et al., 2002; Cuddy et al., 2008; Jeoung et al., 2023), producing a longer list of eleven terms for each axis (Table 1). For surveys of human participants, we use only the Fiske et al. (2002) subset in order to prevent survey fatigue. [^6] These lists were validated at the time of those studies as good operationalizations of their respective concepts, using contemporaneous psychological research methods. The concepts, validation methods, and validation results can change over time. Recent efforts to review the validation of the SCM term lists (Halkias and Diamantopoulos, 2020; Friehs et al., 2022) both affirm and question their value. With this in mind, we re-analyzed data from our rating task (Figure 2) considering only those human and LLM responses that contained words with strong validity according to Halkias and Diamantopoulos (2020). This re-analysis found little change in either the relationship between human and LLM ratings or in overall patterns of results. Hence, we believe that our results will hold under different sets of words that operationalize the concepts of Warmth and Competence. All words in these lists are positive, following Fiske et al. (2002). This is because rating groups on these attributes is part of the survey method, and we confirmed in pilot experiments that LLMs often refused to rate social groups against negative attributes.
@@ 3.2 Models
We tested three representative models that were in widespread use at the time of our survey: [^7] August 2024. GPT 3.5-turbo, Gemini-1.5-flash, and LLaMA 2-7b-chat-hf. GPT and Gemini were accessed by API and LLaMA was run locally. Following pilot experiments, we used a temperature of 0.9 for GPT and Gemini, which were relatively insensitive to this parameter. For LLaMA, we used a temperature of 0.6 and top-$p$ of 0.9; higher temperatures produced output unrelated to the prompt.
The LLMs in our experiments contain safety settings which are intended to block harassment, hate speech, sexual content, and dangerous language. Although our experiments don’t contain such material, they mention social groups that are often targets of such material. In pilot experiments, models frequently refused to produce the requested output. For example, Gemini refused approximately one third of our prompts in both survey and text generation experiments. Analysis showed little difference in output for unblocked prompts, so we turned off safety settings for experiments reported below.
@@
4 SCM Survey of Humans Compared to SCM Prompting of LLMs
Our focus is on LLMs, and we want to understand whether they behave similarly to humans when prompted for stereotypical associations. But the surveys by Fiske et al. (2002) and Cuddy et al. (2007) do not include nonbinary people, lesbians, bisexuals, or heterosexuals. [^8] Note that Jeoung et al. (2023) do include these groups, except for nonbinary people, but their list includes a number of groups not in the key studies that we cite here. In order to compare LLM biases to societal biases towards these groups, we conducted a new survey of human participants. We ask both human participants and LLMs to answer the same types of questions, which probe societal perceptions of groups on the Warmth and Competence axes in three ways:
@@ 4.1 LLM Prompt Design
LLMs were tested using prompts inspired by StereoMap Jeoung et al. (2023), themselves adapted from the survey questions of Fiske et al. (2002), mirroring the three tasks that we gave to human participants. Though Jeoung et al. (2023) did not include a selection task, we include one to better align with previous research on stereotypes (e.g. Cao et al., 2022; Shrawgi et al., 2024; Nadeem et al., 2021). Using multiple question types supports convergent validity of the results, and is more robust to varying rates of refusals.
@@
Table 2: Sample LLM prompts, showing variables in teal. Following StereoMap (Jeoung et al., 2023), the request for models to explain their reasoning, as seen in the first prompt, is based on the rationale of Chain Of Thought (CoT) prompting Wei et al. (2022), which often improves LLM performance on tasks.
@@ 4.2 Associating Elicited Words to SCM Axes
Since elicited words are not necessarily from our predefined lists, we need to know whether any of them represent Warmth and Competence, in order to relate the results to other survey questions. To do this, we use a dictionary created by Nicolas et al. (2021) as a psychological measurement instrument. It associates a large number of terms with several widely-studied stereotype dimensions, partly inspired by the SCM concepts of Warmth and Competence, and has been tested for internal consistency and validity with respect to human judgment, as well as other psychological inventories used to measure these dimensions. For purposes of our analysis, we associate their categories of Morality and Sociability with Warmth, and their categories of Agency and Ability with Competence. Words in these categories account for 45% of the observed word types in our elicited data. For the remaining words, we compute the cosine similarity of their OpenAI word vectors to the average word vector of our Warmth and Competence keywords (Table 1). We assume that those words with a similarity greater than 0.55 represent the associated concept, since this threshold gave us the highest agreement for words in the Nicolas et al. (2021) inventory.
4.3 Human Participants
@@
We recruited participants using the platform Prolific, filtering for English fluency. Ninety-seven participants were each asked to answer twenty-one questions—one of each type, for each of the seven groups in Table 1. They were compensated with an amount above the national minimum wage in the country where we conducted our research. Our survey asked about age range and gender identity, but retained no further identifying information. Most participants were between 16 and 35 (79%), with 14% between 36 and 45, with similar numbers of women and men (54% and 43%, respectively). 6% of participants were over the age of 45 and 2% of participants identified as nonbinary.
4.4 Results
We analyze results below for each task, discussing the human results side by side with LLM results for each task to enable clear comparison.
222233334444222233334444HumanbisexualgayheterosexuallesbianmennonbinarywomenAverage Warmth Rating
3333444433334444GPTbisexualgayheterosexuallesbianmennonbinarywomen
3333444433334444Geminibisexualgayheterosexuallesbianmennonbinarywomen
3333444433334444LLaMAbisexualgayheterosexuallesbianmennonbinarywomen
Average Competence Rating
Figure 2: Average Rating of Warmth and Competence for each group, as given by humans and LLMs. In principle, ratings can range from 1 to 5, but in practice, they fell between 2 and 4 for humans, and between 3 and 4 for LLMs, so we show only those ranges. Note that the LLaMA ratings for nonbinary and bisexual people are nearly identical, so are difficult to distinguish in the visualization.
Rating
@@ 4.4 Results
Figure 2 summarizes results of the rating task. The results clearly show that perceptions of each group do indeed differ according to both humans and LLMs. What is most striking is the pattern for human participants, GPT, and LLaMA: an identical pattern of outliers is clear, with nonbinary people, women, heterosexuals, and men appearing in the same relative positions; and bisexuals, gay men, and lesbians clustered in the center. In all cases, women and men rate most highly in Warmth and Competence, respectively. The other striking result is that the LLM results are shifted towards the more positive end on both axes. The results for women, men, and gay men are consistent with those of Fiske et al. (2002), who did not include the other groups in our suvey. Gemini behaves differently from the other models, with a nearly inverse relationship between Warmth and Competence.
2020-20- 200020202020404040406060606080808080Humanof responses (“not" terms in negative)CompetentNot CompetentWarmNot Warm
00252525255050505075757575100100100100GPTof responses
@@
Figure 3: Percentages of Warmth and Competence terms obtained by Selection. Human percentages include selection of inverse (“not”) terms for each axis. Since this is a selection task, the sum of all percentages is 100% for each group in each graph. To facilitate comparisons of Warmth and Competence, we show them side by side, and to facilitate comparisons of positive and negative terms, we show the negative terms as negative values on the vertical axis. To facilitate comparison with Figure 2, groups are ordered by ascending human rating for Competence. LLaMA is omitted due to high refusal rate.
Selection
We asked participants to select a single term from a list of twenty: a ten-word subset of the Warmth-Competence key terms (Table 1) and an inverse for each positive word, such as ‘cold’ for ‘warm’. Inverse terms were omitted for LLMs, which generally refused to use them. The results (Figure 3) are broadly consistent with the rating results and with previous studies: human participants rate women highest for Warmth; men highest for Competence; nonbinary people most negatively; heterosexuals most positively. They slightly prefer Competence terms for gay men, whereas they slightly preferred Warmth in the rating task. GPT is somewhat consistent with humans, while Gemini tends to skew strongly towards either Warmth or Competence for each group, though this skewed response is internally consistent with its rating results. It prefers Competence for three groups that humans rated most highly for Competence.
@@
Human confused emotional confused flamboyant masculine normal strong